OpenStudy (anonymous):

how does 2cosx(4sinx) become 4sin(2x)

jimthompson5910 (jim_thompson5910):

2*cos(x)*4*sin(x) 2*cos(x)*2*2*sin(x) 2*2*2*cos(x)*sin(x) 4*2*cos(x)*sin(x) 4*( 2*cos(x)*sin(x) ) 4*( sin(2x) ) 4*sin(2x) Note: In step 6, I used the identity 2*sin(x)*cos(x) = sin(2x)

OpenStudy (anonymous):

so pulling out the 2^2 in front legal

jimthompson5910 (jim_thompson5910):

yes I'm using the commutative property of multiplication ie A*B = B*A

OpenStudy (anonymous):

b/c i though about factoring but that wouldn't give me 4sin(2x)

jimthompson5910 (jim_thompson5910):

not sure how you'd factor

OpenStudy (anonymous):

well i jus took out the 2(cosx2sinx)

OpenStudy (anonymous):

then jus did the angle ident

OpenStudy (anonymous):

but i would jus get 2Sin(2x) so i kn am wrong

OpenStudy (anonymous):

is like i lost a power of 2 somewhere